# DP-DyLoRA: Fine-Tuning Transformer-Based Models On-Device under Differentially Private Federated Learning using Dynamic Low-Rank Adaptation

### v3

@@ Abstract
Federated learning (FL) allows clients to collaboratively train a global model without sharing their local data with a server. However, clients’ contributions to the server can still leak sensitive information. Differential privacy (DP) addresses such leakage by providing formal privacy guarantees, with mechanisms that add randomness to the clients’ contributions. The randomness makes it infeasible to train large transformer-based models, common in modern federated learning systems. In this work, we empirically evaluate the practicality of fine-tuning large scale on-device transformer-based models with differential privacy in a federated learning system. We conduct comprehensive experiments on various system properties for tasks spanning a multitude of domains: speech recognition, computer vision (CV) and natural language understanding (NLU). Our results show that full fine-tuning under differentially private federated learning (DP-FL) generally leads to huge performance degradation which can be alleviated by reducing the dimensionality of contributions through parameter-efficient fine-tuning (PEFT). Our benchmarks of existing DP-PEFT methods show that DP-Low-Rank Adaptation (DP-LoRA) consistently outperforms other methods. An even more promising approach, DyLoRA, which makes the low rank variable, when naively combined with FL would straightforwardly break differential privacy. We therefore propose an adaptation method that can be combined with differential privacy and call it DP-DyLoRA. Finally, we are able to reduce the accuracy degradation and word error rate (WER) increase due to DP to less than 2% and 7% respectively with 1 million clients and a stringent privacy budget of $\bm{\epsilon}=2$.
@@ I Introduction
Differential privacy (DP) [7] is the gold standard for providing such privacy guarantees. Very briefly, it adds so much randomness that the data gives very little away about the presence of any individual. Naively applied to the federated learning setting, it would involve adding noise to each individual’s statistics (“local DP”). In this case, the noise would overwhelm the signal. The alternative is to add Gaussian noise once to a sum of many contributions (“central DP”) [8, 9, 10, 11, 12], and use a secure sum algorithm [13, 14, 15] to hide individual contributions from the server.
However, this may still add too much noise. An important lever to change this is the size of the statistics that each client sends to the server [16, 17, 18]. First, if the vector with statistics is longer, its $\ell_{2}$ norm will tend to be greater, and then more noise is required to hide the data. Second, the noise needs to be added to each element of the vector, and therefore the total amount of noise increases with the length of the statistics.
@@
Comprehensive experiments for DP-PEFT methods are missing from the literature. Experiments in existing works [19, 20, 21] fail to address realistic system properties such as a massive number (millions) of clients in a federated learning system [22]. Works such as [20] and [21] show experiments for only a single domain and a single type of DP-PEFT method. Only speech is considered in [19], but it only evaluates Partial Embedding Updates with a combination of LoRA, without considering DP-PEFT methods such as DP-Adapter [23, 24, 25, 26], DP-Compacter [27] and DP-BitFit [28] which are often considered in works regarding PEFT and DP-PEFT methods [29, 30, 31].
This work, on the other hand, presents a comprehensive set of experiments. We start by empirically studying the training dynamics of fine-tuning such models via full fine-tuning on datasets of several domains including natural language understanding, computer vision and speech recognition. We then show with empirical results that parameter-efficient fine-tuning can achieve much better privacy-utility trade-offs than full fine-tuning, and comprehensively benchmark existing DP-PEFT methods on three different domains under DP-FL. The most successful PEFT scheme for DP-FL turns out to be LoRA [29, 30], which learns low-rank matrices to add to existing weight matrices. It is fairly obvious how to use it within DP-FL, where it is called DP-LoRA.
A recent improvement, DyLoRA [32], proposed for NLP, does away with the manual choice of rank. This would make it highly suitable for DP-FL, where privacy budget gets used up by hyperparameter searching. However, a naive adaptation of DyLoRA to DP-FL straightforwardly breaks differential privacy, since users would be sending up vectors of different lengths. In this work, we solve this conundrum by tweaking the form of LoRA so that each cohort uses one vector length. We show that this scheme yields the same expected value of updates as the original DyLoRA. We call the scheme DP-DyLoRA. Our results show that DP-DyLoRA significantly outperforms existing DP-PEFT methods including DP-Adapter [23, 24, 25, 26], DP-Compacter [27], DP-BitFit [28] and DP-LoRA [29] under DP-FL on datasets across three different domains. Specifically, we show that DP-DyLoRA achieves less than 2% accuracy drop and 7% word error rate (WER) increase from non-private LoRA (or DyLoRA) with a strong DP guarantee ($\epsilon=2$) and 1 million clients. DP-DyLoRA achieves noticeably better privacy-utility trade-offs than state-of-the-art DP-PEFT method DP-LoRA as shown in Figure 1.
@@
    Benchmarking existing DP-PEFT methods under DP-FL for a detailed comparison under this learning paradigm.
@@
    Proposing a novel DP-FL algorithm DP-DyLoRA to further improve privacy-utility trade-offs over DP-LoRA by optimising for a range of ranks instead of a fixed rank.
In the remainder of this article, we first present an overview of the related work in Section II, which is followed by preliminaries in Section III. Next, we describe parameter-efficient fine-tuning with differential privacy and introduce our novel DP-FL algorithm DP-DyLoRA in Section IV. We then describe our experimental setup in Section V and discuss our results and findings in Section VI. Finally, we summarise our findings in Section VII.
@@ II Related Work
FederatedAveraging (FedAvg) [33] which served as a generalisation of FederatedSGD (FedSGD) [34] became a common baseline for federated learning soon after being proposed back in 2017. Following the success of FedAvg, numerous works focusing on different aspects of this learning paradigm have been published. As an important research topic for federated learning, various optimisation approaches were proposed to reduce the communication cost and improve robustness against non-independent and identically distributed (non-IID) data [35, 36, 37, 38]. These methods typically attempt to tackle the communication bottleneck of federated learning by either reducing the number of communication rounds required for models to converge or the percentage of clients being sampled at each communication round. Works including [39, 40, 35, 37] focus more on model convergence with large heterogeneity in the data which is a more realistic setting for federated learning [41].
Although federated learning allows clients to contribute to a global model without sharing local data therefore protecting data privacy to some extent, adversaries are still able to infer sensitive information from gradients sent from clients to the server [8, 42, 43], [44]. In order to improve privacy protection in federated learning, DP-FedAvg was proposed in [45] which adds differential privacy to the FedAvg algorithm. This is achieved by introducing noise to the uploaded gradients by using the moments accountant [46] originally proposed for differentially private stochastic gradient descent (DP-SGD) [8] with Gaussian mechanism [47] and privacy amplification via subsampling [9]. The moments accountant provides tight privacy bounds for the sampled Gaussian mechanism [46]. There have also been recent works on training transformer models via DP-SGD which primarily focus on reducing memory and computation complexities [30], [48].
The magnitude of the noise added to achieve differential privacy increases as the model size grows [16, 17, 18]. With the current trend of developing and deploying ever larger models, parameter-efficient fine-tuning turns out to be an intuitive solution for sample-level DP-SGD as proposed in [30] which potentially allows us to fine-tune less than 1% parameters while preserving most of the model performance. Adapter [23, 24, 25, 26] and Low-Rank Adaptation (LoRA) [29] are examples of such methods which can be categorised into sequential and parallel approaches [49], respectively. This approach has been applied also to differentially private federated learning as in [19], [20].
@@ III-A Federated Learning
Federated learning is a machine learning paradigm where a central server aims to train a model on data that is distributed over many clients. What the clients send is not the actual data, but instead statistics about the data. This normally works iteratively: in each round, the server sends the most recent model to a cohort of clients. In Federated Averaging [33], the clients trains for multiple local iterations on local data, and sends the difference between the resulting model and the original one to the server. On the server, the average of updates from all clients turns out to form a good update for the central model, which is the key insight that allows Federated Averaging.
The local training data in a federated learning system is not necessarily independent and identically distributed (IID) [39]. In practice, it is very likely that individual clients train on highly skewed non-IID data [39]. Data heterogeneity can come from different factors such as label distribution and user habit.
@@ III-B Differentially Private Federated Learning
Differential privacy [7] (DP) prevents a membership attack, where an attacker already knows what an individual is sending, but tries to work out whether or not they are included in the data. This seems like a high bar, but no meaningful lower bars have been found. In practice, in federated learning, enough noise must be added to mask any one individual’s statistics. To do this, first the $\ell_{2}$ sensitivity must be constrained, which means making sure that every individual’s contribution $\Delta_{k}$ has $\lVert\Delta_{k}\rVert_{2}<S$, where $S$ is a scalar constant, the clipping bound. Then, noise must be added. The amount of noise is determined ultimately by the privacy budget $(\epsilon,\delta)$, with $\epsilon$ usually being a single-digit value, here $2$, and $\delta$ being a small fraction, here $10^{-6}$.
@@
Assuming such a secure summing algorithm, the differential privacy analysis first proposed in [46] can be used. It assumed a large population of individuals, and at each round of Federated Averaging a subset of them is sampled i.i.d. to contribute. The individual contributions from the selected cohort are summed with added Gaussian noise, and this is used to update the central model. [46] proposed a DP analysis of the algorithm called the “moments accountant”, which was much more efficient in terms of privacy budget than previous analyses, and this analysis has since been improved [9, 10, 11, 12]. In the rest of this article a budget of $(2,10^{-6})$ is used. The Gaussian noise will be chosen to remain within this budget.
Algorithm 1 DP-FedAvg with SecureSum.
1:  Server
2:       parameters
3:           number of communication rounds $T$
4:           total users $\mathcal{K}$
5:           user sampling rate $q\in(0,1]$
6:           noise multiplier $z$
7:           clip norm $S$
8:       initialise model weights $W^{0}$
9:       for each round $t=1,2,\ldots T$ do
10:           Sample a subset $\mathcal{C}^{t}\subseteq\mathcal{K}$ of users uniformly at          random with probability $q$
11:           $\sigma=z\cdot S$
12:           $W^{t+1}\leftarrow W^{t}+\frac{1}{\lvert\mathcal{C}^{t}\rvert}\textsc{SecureSumDP}\big{(}$
13:                $\{\textsc{UserUpdate}(k,W^{t})\}_{k\in\mathcal{C}^{t}},z,S\big{)}$
14:
15:  SecureSumDP($(\{\Delta_{k}\}_{k\in\mathcal{C}^{{}^{\prime}}},z,S)$
16:       parameters
17:            $\sigma=z\cdot S$
18:       return $\mathcal{N}(0,I\sigma^{2})+\sum_{k\in C^{{}^{\prime}}}\Delta_{k}\cdot\min\big{(}1,\frac{S}{\lVert\Delta_{k}\rVert_{2}}\big{)}$
19:
20:  UserUpdate$(W^{{}^{\prime}})$
21:       parameters
22:           number of local epochs $E$
23:           minibatch size $\beta$
24:           learning rate $\eta$
25:       $W^{+}\leftarrow W^{{}^{\prime}}$
26:       for each local epoch $e=1,2,\ldots E$ do
27:           $\mathcal{B}\leftarrow$ (split local data into batches of size $\beta$)
28:           for batch $b\in\mathcal{B}$ do
29:               $W^{+}\leftarrow W^{+}-\eta\triangledown\ell(W^{+})$
30:       return $W^{+}-W^{{}^{\prime}}$
Algorithm 1 shows the algorithm for DP-FedAvg with secure summing. Note that the standard deviation of the noise ($z\cdot S$) is proportional to the clipping bound $S$, which bounds the norm of each individual contribution vector. Also, independent noise is added to each element of the vector. To improve the signal-to-noise ratio, then, and to make training work better, reducing the dimensionality of what each client sends up is a good strategy. First, the vectors will tend to have a lower $\ell_{2}$ norm $\lVert\Delta_{k}\rVert_{2}$, so the clipping bound $S$ can be lowered, and second, the noise is added to fewer elements. The rest of this article will focus on demonstrating the power of this approach.
@@ IV Parameter-Efficient Fine-tuning with Differential Privacy
Parameter-efficient fine-tuning (PEFT) is a technique designed for efficient adaptation of pre-trained models to downstream tasks. Instead of fine-tuning all parameters of a model, parameter-efficient fine-tuning methods aim to train only a small number of parameters. Especially for large pre-trained models, this makes training much cheaper [23, 24, 25, 26, 53, 29].
@@ IV-C BitFit
BitFit is a simple and intuitive parameter-efficient fine-tuning method where only the bias-terms of the pre-trained model are fine-tuned. This method is comprehensively studied in [28] and is often used as a baseline method for PEFT [29], [27].
@@ IV-D LoRA
Low-Rank Adaptation (LoRA) [29] is a parameter-efficient fine-tuning method designed for transformer-based pre-trained models. LoRA can significantly reduce the number of trainable parameters during fine-tuning by freezing the pre-trained weights and adding trainable rank decomposition matrices into each transformer layer. Let $W_{\text{pt}}^{i}\in\mathbb{R}^{b\times a}$ be a pre-trained weight matrix of the $i^{\text{th}}$ layer, LoRA adds a low-rank term $B^{i}A^{i}$ with rank $r$ by:
$W^{i}_{\text{LoRA}}=W_{\text{pt}}^{i}+B^{i}A^{i},$ (3)
where $B^{i}\in\mathbb{R}^{b\times r}$ is an up-projection and $A^{i}\in\mathbb{R}^{r\times a}$ is a down-projection. Here, $A^{i}$ and $B^{i}$ are initialised to random Gaussian noise and zero, respectively. $B^{i}A^{i}$ is therefore zero at the start of training. The pre-trained weights $W_{\text{pt}}^{i}$ are then frozen and the added term $B^{i}A^{i}$ becomes the new trainable parameters.
LoRA has demonstrated superior performance in DP-FL both in central [30] and federated learning [19] settings when compared to other parameter-efficient fine-tuning methods such as adapter [23] and reparametrised gradient perturbation (RGP) [53].
IV-E DyLoRA

### v4

@@ Abstract
Federated learning (FL) allows clients to collaboratively train a global model without sharing their local data with a server. However, clients’ contributions to the server can still leak sensitive information. Differential privacy (DP) addresses such leakage by providing formal privacy guarantees, with mechanisms that add randomness to the clients’ contributions. The randomness makes it infeasible to train large transformer-based models, common in modern federated learning systems. In this work, we empirically evaluate the practicality of fine-tuning large scale on-device transformer-based models with differential privacy in a federated learning system. We conduct comprehensive experiments on various system properties for tasks spanning a multitude of domains: speech recognition, computer vision (CV) and natural language understanding (NLU). Our results show that full fine-tuning under differentially private federated learning (DP-FL) generally leads to huge performance degradation which can be alleviated by reducing the dimensionality of contributions through parameter-efficient fine-tuning (PEFT). Our benchmarks of existing DP-PEFT methods show that DP-Low-Rank Adaptation (DP-LoRA) and its variants consistently outperform other methods. An even more promising approach, DyLoRA, which makes the low rank variable, when naively combined with FL would straightforwardly break differential privacy. We therefore propose an adaptation method that can be combined with differential privacy and call it DP-DyLoRA. Finally, we are able to reduce the accuracy degradation and word error rate (WER) increase due to DP to less than 2% and 7% respectively with 1 million clients and a stringent privacy budget of $\bm{\epsilon}=2$.
@@ I Introduction
Differential privacy (DP) [7] is the gold standard for providing such privacy guarantees. Very briefly, it adds so much randomness that the data gives very little away about the presence of any individual. Naively applied to the federated learning setting, it would involve adding noise to each individual’s statistics (“local DP”). In this case, the noise would overwhelm the signal. The alternative is to add Gaussian noise once to a sum of many contributions (“central DP”) [8, 9, 10], and use a secure sum algorithm [11, 12, 13] to hide individual contributions from the server.
However, this may still add too much noise. An important lever to change this is the size of the statistics that each client sends to the server [14, 15]. First, if the vector with statistics is longer, its $\ell_{2}$ norm will tend to be greater, and then more noise is required to hide the data. Second, the noise needs to be added to each element of the vector, and therefore the total amount of noise increases with the length of the statistics.
@@
Comprehensive experiments for DP-PEFT methods are missing from the literature. Experiments in existing works [16, 17, 18] fail to address realistic system properties such as a massive number (millions) of clients in a federated learning system [19]. Works such as [17] and [18] show experiments for only a single domain and a single type of DP-PEFT method. [16] only considers the speech domain and evaluates Partial Embedding Updates with a combination of LoRA, without considering PEFT methods such as Adapter [20, 21], Compacter [22] and BitFit [23] which are often considered in works regarding PEFT and DP-PEFT methods [24, 25, 26].
This work, on the other hand, presents a comprehensive set of experiments. We start by empirically studying the training dynamics of fine-tuning transformer-based models via full fine-tuning on datasets of multiple domains including natural language understanding, computer vision and speech recognition. We then show with empirical results that parameter-efficient fine-tuning can achieve much better privacy-utility trade-offs than full fine-tuning, and comprehensively benchmark existing DP-PEFT methods on three different domains under DP-FL. The most successful PEFT scheme for DP-FL turns out to be LoRA [24, 25], which learns low-rank matrices to add to existing weight matrices. It is fairly obvious how to use it within DP-FL, where it is called DP-LoRA.
A recent improvement, DyLoRA [27], proposed for NLP, does away with the manual choice of rank. This would make it highly suitable for DP-FL, where privacy budget gets used up by hyperparameter searching. However, a naive adaptation of DyLoRA to DP-FL straightforwardly breaks differential privacy, since users would be sending up vectors of different lengths. In this work, we solve this conundrum by modifying the form of DyLoRA so that each cohort uses one vector length. We show that this scheme yields the same expected value of updates as the original DyLoRA. We call the scheme DP-DyLoRA. Our results show that DP-DyLoRA significantly outperforms existing DP-PEFT methods including DP-Adapter [20, 21], DP-Compacter [22], DP-BitFit [23], DP-LoRA [24] and its variants DP-LoHa [28] and DP-AdaLoRA [29] under DP-FL on datasets across three different domains. Specifically, we show that DP-DyLoRA achieves less than 2% accuracy drop and 7% word error rate (WER) increase from non-private LoRA (or DyLoRA) with a strong DP guarantee ($\epsilon=2$) and 1 million clients. DP-DyLoRA achieves noticeably better privacy-utility trade-offs than state-of-the-art DP-PEFT method DP-LoRA as shown in Figure 1.
@@
    First work in literature to comprehensively benchmark existing PEFT methods under DP-FL with production-level system parameters for a detailed and realistic comparison under this learning paradigm.
@@
    Proposing a novel DP-FL algorithm DP-DyLoRA which achieves better privacy-utility trade-offs than state-of-the-art DP-PEFT method DP-LoRA by optimising for a range of ranks instead of a fixed rank.
  3. 3.
    Formal security analysis to prove that our proposed algorithm satisfies differential privacy constraints.
The rest of the paper is organised as follows. We first present an overview of the related work in Section II, which is followed by preliminaries in Section III. Next, we describe parameter-efficient fine-tuning with differential privacy in Section IV and introduce our novel DP-FL algorithm DP-DyLoRA with formal security analysis in Section V. We then describe our experimental setup in Section VI and discuss our results and findings in Section VII. Finally, we summarise our findings in Section VIII.
@@ II Related Work
FederatedAveraging (FedAvg) [30] which served as a generalisation of FederatedSGD (FedSGD) [31] became a common baseline for federated learning (FL) soon after being proposed back in 2017. Following the success of FedAvg, numerous works focusing on different aspects of this learning paradigm have been published [32, 33, 34]. As an important research topic for FL, various optimisation approaches were proposed to reduce the communication cost and improve robustness against non-independent and identically distributed (non-IID) data [32, 33, 35, 34]. These methods typically attempt to tackle the communication bottleneck of FL by either reducing the number of communication rounds required for models to converge or the percentage of clients being sampled at each communication round. Works including [36, 37, 32, 35] focus more on model convergence with large heterogeneity in the data which is a more realistic setting for FL [38].
Although federated learning allows clients to contribute to a global model without sharing local data therefore protecting data privacy to some extent, adversaries are still able to infer sensitive information from gradients sent from clients to the server [8, 39, 40, 41]. In order to improve privacy protection in federated learning, DP-FedAvg was proposed in [42] which adds differential privacy to the FedAvg algorithm. This is achieved by introducing noise to the uploaded gradients by using the moments accountant [43] originally proposed for differentially private stochastic gradient descent (DP-SGD) [8] with Gaussian mechanism [44] and privacy amplification via subsampling [9]. The moments accountant provides tight privacy bounds for the sampled Gaussian mechanism [43]. There have also been recent works on training transformer models via DP-SGD which primarily focus on reducing memory and computation complexities [25, 45].
The magnitude of the noise added to achieve differential privacy increases as the model size grows [14, 15]. With the current trend of developing and deploying ever larger models, parameter-efficient fine-tuning turns out to be an intuitive solution for sample-level DP-SGD as proposed in [25] which potentially allows us to fine-tune less than 1% parameters while preserving most of the model performance. Adapter [20, 21] and Low-Rank Adaptation (LoRA) [24] are examples of such methods which can be categorised into sequential and parallel approaches [46], respectively. This approach has been applied also to differentially private federated learning as in [16, 17].
@@ III-A Federated Learning
Federated learning is a machine learning paradigm where a central server aims to train a model on data that is distributed over a large number of clients. What the clients send is not the actual data, but instead statistics about the data. This normally works iteratively: in each round, the server sends the most recent model to a cohort of clients. In Federated Averaging [30], the clients trains for multiple local iterations on local data, and sends the difference between the resulting model and the original one to the server. On the server, the average of updates from all clients turns out to form a good update for the central model, which is the key insight that allows Federated Averaging.
The local training data in a federated learning system is not necessarily independent and identically distributed (IID) [36]. In practice, it is very likely that individual clients train on highly skewed non-IID data [36]. Data heterogeneity can come from different factors such as label distribution, number of samples per client and user habit.
@@ III-B Differentially Private Federated Learning
Differential privacy [7] (DP) prevents a membership attack, where an attacker already knows what an individual is sending, but tries to work out whether they are included in the data. This seems like a high bar, but no meaningful lower bars have been found. In practice, in federated learning, enough noise must be added to mask any one individual’s statistics. To do this, first the $\ell_{2}$ sensitivity must be constrained, which means making sure that every individual’s contribution $\Delta_{k}$ has $\lVert\Delta_{k}\rVert_{2}<S$, where $S$ is a scalar constant, the clipping bound. Then, noise must be added. The amount of noise is determined ultimately by the privacy budget $(\epsilon,\delta)$, with $\epsilon$ usually being a single-digit value, here $2$, and $\delta$ being a small fraction, here $10^{-6}$.
@@
Assuming such a secure summing algorithm, the differential privacy analysis first proposed in [43] can be used. It assumed a large population of individuals, and at each round of Federated Averaging a subset of them is sampled i.i.d. to contribute. The individual contributions from the selected cohort are summed with added Gaussian noise, and this is used to update the central model. [43] proposed a DP analysis of the algorithm called the “moments accountant”, which was much more efficient in terms of privacy budget than previous analyses, and this analysis has since been improved [9, 10]. In the rest of this article a budget of $(2,10^{-6})$ is used. The Gaussian noise will be chosen to remain within this budget.
@@ IV Parameter-Efficient Fine-tuning with Differential Privacy
Parameter-efficient fine-tuning (PEFT) is a technique designed for efficient adaptation of pre-trained models to downstream tasks. Instead of fine-tuning all parameters of a model, parameter-efficient fine-tuning methods aim to train only a small number of parameters. This makes training much cheaper especially for large pre-trained models [20, 21, 24].
@@
Here, we describe state-of-the-art PEFT methods which we include later in our benchmark. Most of these methods have been studied under DP constraints such as DP-Adapter, DP-Compacter and DP-LoRA in [25]. All of the PEFT methods we describe in this section can be directly applied to the DP-FedAvg [42] algorithm without modifying either the algorithm or the PEFT method.
@@ IV-C BitFit
BitFit is a simple and intuitive parameter-efficient fine-tuning method where only the bias terms of the pre-trained model are fine-tuned. This method is comprehensively studied in [23] and is often used as a baseline method for PEFT studies [22, 24].
@@ IV-D LoRA
Low-Rank Adaptation (LoRA) [24] is a parameter-efficient fine-tuning method designed for transformer-based pre-trained models. LoRA can significantly reduce the number of trainable parameters during fine-tuning by freezing the pre-trained weights and adding trainable rank decomposition matrices into each transformer layer. Let $W_{\text{pt}}^{i}\in\mathbb{R}^{b\times a}$ be pre-trained weight matrices of the $i^{\text{th}}$ layer, LoRA adds a low-rank term $B^{i}A^{i}$ with rank $r$ by:
$W^{i}_{\text{LoRA}}=W_{\text{pt}}^{i}+B^{i}A^{i},\vspace*{-4pt}$ (3)
where $B^{i}\in\mathbb{R}^{b\times r}$ is an up-projection and $A^{i}\in\mathbb{R}^{r\times a}$ is a down-projection. Here, $A^{i}$ and $B^{i}$ are initialised to random Gaussian noise and zero, respectively. $B^{i}A^{i}$ is therefore zero at the start of training. The pre-trained weights $W_{\text{pt}}^{i}$ are then frozen and $B^{i}A^{i}$ becomes the new trainable parameters.
LoRA has demonstrated superior performance in DP-FL both in central [25] and federated learning [16] settings when compared to other parameter-efficient fine-tuning methods such as adapter [20] and compacter [22].
IV-E LoHa
Low-Rank Hadamard Product (LoHa) or FedPara [28] is a parameter-efficient fine-tuning method proposed specifically for reducing the communication costs in federated learning. Unlike Lora, LoHa uses Hadamard product to approximate larger weight matrices as:
$W^{i}_{\text{LoHa}}=W_{\text{pt}}^{i}+(B_{1}^{i}A_{1}^{i})\odot(B_{2}^{i}A_{2}^{i}),\vspace*{-4pt}$ (4)
where $\odot$ denotes the Hadamard product.
IV-F AdaLoRA
Adaptive Low-Rank Adaptation (AdaLoRA) [29] adopts singular value pruning to adaptively optimise the rank values for different weight matrices based on the magnitude of individual singular values. The weight matrices $W^{i}_{\text{AdaLoRA}}$ for the $i^{\text{th}}$ layer is then defined as
$W^{i}_{\text{AdaLoRA}}=W_{\text{pt}}^{i}+B^{i}\Lambda^{i}A^{i},\vspace*{-4pt}$ (5)
where $\Lambda^{i}\in\mathbb{R}^{r\times r}$ denotes the singular values.
IV-G DyLoRA